OpenStudy (anonymous):

A bank runs a contest to encourage new customers to open accounts. In the contest, each contestant draws a slip representing a different reward—$5, $3, or $x—from a jar. At the beginning of the contest the jar contains 60 slips for $5, 40 slips for $3, and 50 slips for $x. If the expected value of the first draw from the jar is $5.8, the value of x is . At one point in the contest, the jar contains 3 slips for $5, 7 slips for $3, and y slips for $x. If the expected value on the next draw is $6, the value of y is

OpenStudy (anonymous):

Help Please

OpenStudy (dogzcatz):

Well The amount would be 15+21.

OpenStudy (anonymous):

For the value of x? What about the value of y?

OpenStudy (dogzcatz):

I believe the 2nd one would be 12

OpenStudy (dogzcatz):

do you have any Ideas for the 1st one?

OpenStudy (anonymous):

Nope I've been stuck on this problem all day

OpenStudy (anonymous):

The first one wasn't 15+21?

OpenStudy (dogzcatz):

that was for the 2nd one

OpenStudy (anonymous):

Do you know how to do the first one?

OpenStudy (dogzcatz):

156 is the total amount of dollars in the x pot

OpenStudy (anonymous):

So 156=x and 12=y

OpenStudy (anonymous):

Think that's wrong :(

OpenStudy (kropot72):

Let the reward be Y.  \[\large E(Y)=5.8=5\times\frac{60}{150}+3\times\frac{40}{150}+\frac{50x}{150}\ .......(1)\] Next you need to solve equation (1) to find the value of x.

OpenStudy (kropot72):

Equation (1) simplifies to: \[\large 5.8=2+0.8+\frac{x}{3}\]

OpenStudy (anonymous):

I do that to get X? How do I get Y?

OpenStudy (kropot72):

Once you have calculated the value of x, you can set up a probability distribution table similar to the table I have shown you. In the second case the total number of slips is: 3 + 7 + y so the probability of $5 is given by: \[\large P(Y=5)=\frac{3}{3+7+y}\] the probability of $3 is given by: \[\large P(Y=3)=\frac{7}{3+7+y}\] and the probability of $x (put in the calculated value of x here) is given by: \[\large P(Y=x)=\frac{y}{3+7+y}\]

OpenStudy (kropot72):

Have you calculated the value of x yet?